Find the integrating factor of given differential equation

(x2y + y2) dx + (y3-x3) dy = 0


1
Expert's answer
2021-02-24T07:51:04-0500

This Given Differential Equation is of the form:

"M(x,y)dx+N(x,y)dy=0"


where "M(x,y)=x^2y+y^2" and "N(x,y)=y^3\u2212x^3" Now


"Mx\u2212Ny=(x^2y+y^2)x-(y^3-x^3)y"

"\\=x^3y+xy^2-y^4+x^3y\\=2x^3y+xy^2-y^4\\≠0"


Therefore the integrating factor "( I.F.)=\dfrac{1}{Mx-Ny}"


"=\dfrac{1}{2x^3y+xy^2-y^4}"